Trios:

So I’ve recently moved out of equations. I still don’t understand it quite yet. Expample: 8x+4(-6x+4)=-6-5x This isn’t homework or anything just a practice equation. Can someone answer and explain how to get the answer?

Ruthy:

Open bracket

Trios:

Mhm 🤔

darkknight:

8x+4(-6x+4)=-6-5x distribute 8x-24x+16 = -6-5x -16x+ 16 = -6-5x -11x = -22 x = ?

darkknight:

can you follow along? I distributed, then combined like terms on the right and left, then isolated and solved for x

Trios:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @darkknight can you follow along? I distributed, then combined like terms on the right and left, then isolated and solved for x \(\color{#0cbb34}{\text{End of Quote}}\) Ohh then divide 22 by both sides

Trios:

I mean -11

Ruthy:

Mn\(\color{#0cbb34}{\text{Originally Posted by}}\) @Trios \(\color{#0cbb34}{\text{Originally Posted by}}\) @darkknight can you follow along? I distributed, then combined like terms on the right and left, then isolated and solved for x \(\color{#0cbb34}{\text{End of Quote}}\) Ohh then divide 22 by both sides \(\color{#0cbb34}{\text{End of Quote}noooo divide both sides by-11

darkknight:

yep correct @trios

Ruthy:

Then ur ans is 2

Trios:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Ruthy Mn\(\color{#0cbb34}{\text{Originally Posted by}}\) @Trios \(\color{#0cbb34}{\text{Originally Posted by}}\) @darkknight can you follow along? I distributed, then combined like terms on the right and left, then isolated and solved for x \(\color{#0cbb34}{\text{End of Quote}}\) Ohh then divide 22 by both sides \(\color{#0cbb34}{\text{End of Quote}noooo divide both sides by-11 \(\color{#0cbb34}{\text{End of Quote}}\) Yeah lol sorry

Trios:

Thank you guys so much

Ruthy:

No probs u re welcome

darkknight:

np

Ruthy:

Any time u need help we re always here